# Gradient corrections to the local-equilibrium energy-momentum tensor in the Zubarev approach

### v2

@@ III Pseudogauge dependence and gradient-corrected space-time dependent energy-momentum tensor
One observes that $\epsilon_{\text{gc-$z$-geq}}$ is the (space-time dependent) energy density and $u^{\mu}$ is the hydrodynamical four-velocity in the Landau frame of $\tilde{\mathcal{T}}^{\mu\nu}_{\text{gc-$z$-geq}}(z)$. Using Eqs. (47), (48), (49), and (50) we conclude that $\tilde{\mathcal{T}}^{\mu\nu}_{\text{gc-$z$-geq}}(z)$ leads to the differential equations of ideal hydrodynamics with an equation of state, which, for $\Lambda\neq 0$, is not equal to that in (space-time dependent) global thermodynamical equilibrium.
@@
It is convenient to introduce the following notations
@@
For the purpose of illustration, let us consider the non-relativistic, $m\beta=m/T\gg 1$, and ultrarelativistic, $m\beta=m/T\ll 1$, limits. Equation (57) then becomes
@@
where $\lambda_{\text{th}}$ is the thermal wavelength: $\lambda_{\text{th}}\sim 1/\sqrt{mT}$ for $m\beta\gg 1$ and $\lambda_{\text{th}}\sim 1/T$ for $m\beta\ll 1$. Meanwhile, $L$ can be interpreted as the space-time length of homogeneity of a system, see Eq. (54). We thus conclude that the corrections to the global-equilibrium equation of state can be neglected if the thermal wavelengths are much smaller than the characteristic space-time scales of the system,
@@
Otherwise, if $\lambda_{\text{th}}/L\gtrsim 1$, then the energy density, $\epsilon_{z-\text{leq}}$, or pressure, ${\cal P}_{z-\text{leq}}$, can become negative, see Eqs. (48), (49), and (50). Such a violation of the positivity condition implies the inapplicability of hydrodynamics, [^9] See also Ref. small , where the domain of validity of hydrodynamics applied to small systems was considered. and suggests that space-time gradients in the system are so large that one can neither neglect the standard dissipative corrections, nor those arising from space-time gradients of the ideal-hydrodynamical variables such as $\beta_{\nu}$. It is worth noting that the inequality (59) is reminiscent of the condition that the Knudsen number should be sufficiently small for the applicability of the hydrodynamical description. Then, our analysis supports the view that, as an effective theory, hydrodynamics is applicable if there is a clear separation between the microscopic and the macroscopic scales.
@@ IV Conclusions
In the present paper, we calculated non-dissipative corrections to the energy-momentum tensor of a real scalar field in a state which is described by a statistical operator with a space-time dependent temperature four-vector and which is commonly identified with a local-equilibrium state. In particular, we demonstrated that the relation between pressure and energy density, i.e., the equation of state, is affected by these corrections and therefore is modified with respect to the case of global thermodynamical equilibrium. The corrections are of second order in the ratio of the thermal wavelength to the typical macroscopic length-scale of the system. If the thermal wavelengths are comparable in size with the typical spatio-temporal scales of the system, then a hydrodynamical description might fail. Our findings support the conjecture that hydrodynamics is based on a clear separation of scales: the microscopic scales, which are characteristic of the underlying microscopic theory, e.g., a quantum field theory, and the macroscopic scales, which characterize the spatio-temporal variation of the hydrodynamical variables. Our results present a challenge for the hydrodynamical description of the small, inhomogeneous systems created in hadron or hadron-ion collisions.

### v3

@@ III Pseudogauge dependence and gradient-corrected spacetime dependent energy-momentum tensor
One observes that $\epsilon_{\text{gc-$z$-geq}}$ is the (spacetime dependent) energy density and $u^{\mu}$ is the hydrodynamical four-velocity in the Landau frame of $\tilde{\mathcal{T}}^{\mu\nu}_{\text{gc-$z$-geq}}(z)$. Using Eqs. (47), (48), (49), and (50) we conclude that $\tilde{\mathcal{T}}^{\mu\nu}_{\text{gc-$z$-geq}}(z)$ leads to the differential equations of ideal hydrodynamics with an equation of state, which, for $\Lambda\neq 0$, is not equal to that in (spacetime dependent) global thermodynamical equilibrium.
@@
It is convenient to introduce the following notations:
@@
For the purpose of illustration, let us consider the nonrelativistic, $m\beta=m/T\gg 1$, and ultrarelativistic, $m\beta=m/T\ll 1$, limits. Equation (57) then becomes
@@
where $\lambda_{\text{th}}$ is the thermal wavelength: $\lambda_{\text{th}}\sim 1/\sqrt{mT}$ for $m\beta\gg 1$ and $\lambda_{\text{th}}\sim 1/T$ for $m\beta\ll 1$. Meanwhile, $L$ can be interpreted as the spacetime length of homogeneity of a system, see Eq. (54). We thus conclude that the corrections to the global-equilibrium equation of state can be neglected if the thermal wavelengths are much smaller than the characteristic spacetime scales of the system,
@@
Otherwise, if $\lambda_{\text{th}}/L\gtrsim 1$, then the energy density, $\epsilon_{z-\text{leq}}$, or pressure, ${\cal P}_{z-\text{leq}}$, can become negative, see Eqs. (48), (49), and (50). Such a violation of the positivity condition implies the inapplicability of hydrodynamics, [^9] See also Ref. small , where the domain of validity of hydrodynamics applied to small systems was considered. and suggests that spacetime gradients in the system are so large that one can neither neglect the standard dissipative corrections, nor those arising from spacetime gradients of the ideal hydrodynamical variables such as $\beta_{\nu}$. It is worth noting that the inequality (59) is reminiscent of the condition that the Knudsen number should be sufficiently small for the applicability of the hydrodynamical description. Then, our analysis supports the view that, as an effective theory, hydrodynamics is applicable if there is a clear separation between the microscopic and the macroscopic scales.
@@ IV Conclusions
In the present paper, we calculated nondissipative corrections to the energy-momentum tensor of a real scalar field in a state which is described by a statistical operator with a spacetime dependent temperature four-vector and which is commonly identified with a local-equilibrium state. In particular, we demonstrated that the relation between pressure and energy density, i.e., the equation of state, is affected by these corrections and therefore is modified with respect to the case of global thermodynamical equilibrium. The corrections are of second order in the ratio of the thermal wavelength to the typical macroscopic length scale of the system. If the thermal wavelengths are comparable in size with the typical spatiotemporal scales of the system, then a hydrodynamical description might fail. Our findings support the conjecture that hydrodynamics is based on a clear separation of scales: the microscopic scales, which are characteristic of the underlying microscopic theory, e.g., a quantum field theory, and the macroscopic scales, which characterize the spatiotemporal variation of the hydrodynamical variables. Our results present a challenge for the hydrodynamical description of the small, inhomogeneous systems created in hadron or hadron-ion collisions.